7. What is meant by the equilibrium-relative commod￾ity price in isolation? How is this price determined


in each nation? How does it define the nation’s


comparative advantage?

1
Expert's answer
2022-01-23T15:39:45-0500

In absence of trade, the relative commodity price at which the country should maximize the welfare. It is the price where the country has been maximized the welfare of the people. It is determined by the slope of tangent between production possibility frontier (PPF) and indifference curve at the isolation point of production of a nation and there is summation indifference curve of the whole nation.

As we know relative commodity price in isolation means that comparison between two goods prices of the same nation. Its equilibrium has been determined by two different variables such as the production and demand of the different commodities.

b). The price has been determined by the slope of the production i.e supply whereas the nation's indifference curve i.e., demand price of each nation have been different it can be determined by the relative price of two commodities.

Example:

The price of Nation 1 can be determined by

P1= Price of Good X / Price of Good Y

The price of Nation 2 can be determined by

P2= Price of Good X / Price of Good Y

Here, the price of both nations is different in the case of isolation whereas when there is trade adjusted price has been identified so no one can harm from it.

c). As we know comparative advantage means specialization or less cost incurred in producing something as compared to other nations.

Example:

Suppose that price of the commodity Good Y in Nation 1 is $1 and Nation 2 is $6, whereas the price of Commodity Good Z in Nation 1 is $6 and Nation 2 is $1.

P1= PY/ Pz= 1/6where as P2= PY/ PZ= 6/1

P1< P2

So, Nation 1 has a comparative advantage in commodity Y whereas Nation 2 has a comparative advantage in commodity Z.